OpenStudy (anonymous):

Find n so that the number of nP2 is 6 times as great as the number of permutations of n-5 elements taken 2 at a time.

OpenStudy (amistre64):

maybe .... n(n-1) = 6(n-5)(n-6)

OpenStudy (amistre64):

got me 6 on the worng side i think

OpenStudy (anonymous):

\[_{n} P_{2} = 6\left( _{n-5}P_{2} \right)\] \[\frac{ n! }{ \left( n-2 \right)! }=6 \left( \frac {(n-5)!} {[(n-5)-2]!} \right)\] \[\frac{ n! }{ \left( n-2 \right)! }=6 \left( \frac {(n-5)!} {(n-7)!} \right)\] \[n (n-1)=6 (n-5)(n-6)\] \[n^2-n=6 (n^2-11n+30)\] \[n^2-n=6n^2-66n+180\] \[0=5n^2-65n+180\] \[0=5(n^2-13n+36)\] \[0=5(n-4)(n-9)\] \[n = 4\] or \[n=9\]. Since \[n \neq 4\] because \[ n \ge 7\] then \[ n= 9\]

OpenStudy (amistre64):

yeah, the 6 on the right is gooder

OpenStudy (anonymous):

Check. \[ _{9}P_{7} = 6(_{4}P_{2})\] \[9(8) = 6(4)(3) \] \[72 = 72 \]